# AMMeBa: A Large-Scale Survey and Dataset of Media-Based Misinformation In-The-Wild

### v1

@@ 8 Conclusion
At first blush, misinformation is often conceptualized in a sense as essentially incorrect statements about the world which may predispose us to think of it as being rendered either verbally or through text. Recently, though, it’s become clear that misinformation often relies on the use of media, like images or video, to articulate or support a false claim. Indeed “often” is perhaps an understatement, as we show that about 80% of fact-checked misinformation claims involve media of some kind in a material way. Research efforts must be cognizant of this, and be conducted with this multi-modality in mind. Images and video occur regularly, with video-based misinfo becoming most common modality recently, possibly due to the rising popularity of video platforms.
Pixel-based forensic methods are a commonly-cited means of identifying image-based manipulations. These methods typically identify content manipulations through irregularities in low-level image properties, like compression and noise artifacts. They are the subject of considerable research, particularly as generative AI becomes an increasing concern. However, we find image manipulations historically tended to be simple. In particular, context manipulations—where images are paired with false claims about what they depict—are extremely common. Context manipulations often use authentic images and involve no deceptive content manipulations at all, making pixel-based forensic methods ineffective in those cases. This points to a need for approaches that can flexibly incorporate contextual information, and compare it to other versions of the media available online. Identifying such manipulations in general is an open research problem, though it’s likely some categories, particularly Date/Time and Location, can by automatically inferred given other documents bearing the image.
@@
Context manipulations are not always explicit. Frequently, they’re made by implication (or by functioning as “dog whistles”), not only manipulating the context of the image but also requiring context of their own to apprehend their meaning. These manipulations may also rely on the interaction of multiple sub-images alongside text, including text in the image itself (examples of this are shown in Fig. 27). This complexity underscores an a crucial fact: the simplicity of a manipulation does not imply simplicity of detection.
AI-generated content is a prominent topic in discussions of online misinformation. Our data show that, historically, this was not reflected in the online population of misinformation claims, as AI-generated content was extremely rare, at least among claims that are fact-checked. This changed abruptly in 2023, where AI-generated images began to rise rapidly as a fraction of overall rated images, roughly in line with the viral popularity of the “puffy jacket” image of the Pope. While AI-generated images did not cause content manipulations to overtake context manipulations, our data collection ended in late 2023 and this may have changed since. Regardless, generative-AI images are now a sizeable fraction of all misinformation-associated images.
@@ 9.3 Data Attrition
Because fact checks were assessed in multiple stages in our study at different points in time, we could estimate the rate at which it occurs by finding cases where the link identified by the rater in one stage fails to work in subsequent stage. We find attrition occurs most rapidly in the time immediately after the fact check is published, followed by a slow decay thereafter, which gradually accelerates over time.
Attrition is a limitation of this study, as we it’s unlikely it occurs in a random manner; we expect more virulent or harmful misinformation claims are more likely to undergo attrition than relatively benign examples. This phenomenon is not only a problem for this study, but also in general, as it diminishes the utility of the fact check if it prevents those accessing it from knowing what claims it actually addresses. Many fact checkers use archival sites to overcome this problem, but their use is not universal and their reliability in general is not known.
@@ 9.5 Finer Categorization
In order to properly study the breadth of context manipulations, subsequent stages would be necessary where raters could apply finer-grained annotations. Further deep dives, particularly in categories that exhibit large within-category variation or with large numbers of examples that defy subcategorization would afford greater understanding.
@@ Appendix A Inclusion Criteria
Lastly, fact checks that did not regard the claim as false, misleading or misinformation were also not further annotated. Fact checking organizations use a variety of heterogeneous labelling systems to indicate veracity, raters were instructed to err on the side of excluding misinformation claims that were assessed as “possibly true” or “unclear,” etc.
@@ B.1 Image Types
The image type categories were selected to reflect properties of the images that are relevant to an investigator, like a fact checker, engaged in determining its authenticity, as the means used for investigating an image present in a misinformation claim differ based on the nature of the image and the nature of the suspected manipulation. For example, while reverse image search methods are reliable in the case of natural images and are widely deployed in fact checking [18], they may be less useful for images with large amounts of diagrammatic content and text [16], which contain less information than natural images and bear elements that are frequently and identically repeated across images. Additionally, content may be captured by taking a screenshot and re-shared, sometimes several times. Each of these can introduce graphical elements into the new image, as well as reduce the fidelity of the underlying original.
@@ B.3.1 Fake Official Documents
Fake official documents are create to give the false impression of an official communication from an organization or authority. This category was not present in our original formulation of the typologies, but was added once their prevalence was recognized. These types of images are particularly relevant during public safety events, where communication from officials can be crucial.
@@ Appendix C Data Collection Rater Task
In all cases, raters were instructed to treat the fact check as a source of truth for annotations, and to refrain from making independent judgements. Exceptions to this are cases where the judgement could be made objectively or concerned categories developed specifically for this study (principally, the distinction between basic and complex images, see 5.3) and never when related to the veracity of the misinformation claim being examined or the evidence establishing this veracity. In cases where misinformation claims are addressed by multiple fact checks presenting conflicting assessments of claims, or raters provided conflicting responses, those conflicts will be reflected in the dataset.
@@ C.3 Stage 2: Fine-Grained Annotation of Single-Image Misinformation Claims
Stage 2 contains dozens of categories, most centering on the manipulation present in, or applied to, the image in order to support or make the misinformation claim (does the image contain AI-based manipulation, did the fact check conduct a reverse-image search, etc). In contrast with Stages 1 and 1M, very few annotations were discarded (a total of 69). Each (claim, image) pair has an average of 2.6 Stage 2 annotations, and 84.9% of claims have more than 1 replication.
@@ C.4 Stage 2M: Fine-Grained Annotation of Images from Multiple-Media Misinformation Claims
Stage 2M was the final stage conducted, and is applied to misinformation claims bearing multiple media including at least one image. For every image associated with such a claim, a Stage 2M annotation is applied. Raters are presented with a fact check URL and an original source as identified in Stage 1. Images associated with the claim, identified in Stage 1M (see Sec. C.2), are also shown; one image is the "primary" image, with the remainder secondary. The primary image was is the main subject of annotation.

### v2

@@ 8 Conclusion
At first blush, misinformation is often conceptualized in a sense as essentially incorrect statements about the world which may predispose us to think of it as being rendered either verbally or through text. Recently, though, it’s become clear that misinformation often relies on the use of media, like images or video, to articulate or support a false claim. Indeed “often” is perhaps an understatement, as we show that about 80% of fact-checked misinformation claims involve media of some kind in a material way. Research efforts must be cognizant of this, and be conducted with this multi-modality in mind. Images and video occur regularly, with video-based misinfo becoming the most common modality recently, possibly due to the rising popularity of video platforms.
Pixel-based forensic methods are a commonly-cited means of identifying image-based manipulations. These methods typically identify content manipulations through irregularities in low-level image properties, like compression and noise artifacts. They are the subject of considerable research, particularly as generative AI becomes an increasing concern. However, we find image manipulations historically tended to be simple. In particular, context manipulations—where images are paired with false claims about what they depict—are extremely common. Context manipulations often use authentic images and involve no deceptive content manipulations at all, making pixel-based forensic methods ineffective in those cases. This points to a need for approaches that can flexibly incorporate contextual information, and compare it to other versions of the media available online. Identifying such manipulations in general is an open research problem, though it’s likely some categories, particularly Date/Time and Location, can be automatically inferred given other documents bearing the image.
@@
Context manipulations are not always explicit. Frequently, they’re made by implication (or by functioning as “dog whistles”), not only manipulating the context of the image but also requiring context of their own to apprehend their meaning. These manipulations may also rely on the interaction of multiple sub-images alongside text, including text in the image itself (examples of this are shown in Fig. 27). This complexity underscores a crucial fact: the simplicity of a manipulation does not imply simplicity of detection.
AI-generated content is a prominent topic in discussions of online misinformation. Our data show that, historically, this was not reflected in the online population of misinformation claims, as AI-generated content was extremely rare, at least among claims that are fact-checked. This changed abruptly in 2023, where AI-generated images began to rise rapidly as a fraction of overall rated images, roughly in line with the viral popularity of the “puffy jacket” image of the Pope. While AI-generated images did not cause content manipulations to overtake context manipulations, our data collection ended in late 2023 and this may have changed since. Regardless, generative-AI images are now a sizable fraction of all misinformation-associated images.
@@ 9.3 Data Attrition
Because fact checks were assessed in multiple stages in our study at different points in time, we could estimate the rate at which it occurs by finding cases where the link identified by the rater in one stage fails to work in subsequent stages. We find attrition occurs most rapidly in the time immediately after the fact check is published, followed by a slow decay thereafter, which gradually accelerates over time.
Attrition is a limitation of this study, as it’s unlikely it occurs in a random manner; we expect more virulent or harmful misinformation claims are more likely to undergo attrition than relatively benign examples. This phenomenon is not only a problem for this study, but also in general, as it diminishes the utility of the fact check if it prevents those accessing it from knowing what claims it actually addresses. Many fact checkers use archival sites to overcome this problem, but their use is not universal and their reliability in general is not known.
@@ 9.5 Finer Categorization
To properly study the breadth of context manipulations, subsequent stages would be necessary where raters could apply finer-grained annotations. Further deep dives, particularly in categories that exhibit large within-category variation or with large numbers of examples that defy subcategorization would afford greater understanding.
@@ Appendix A Inclusion Criteria
Lastly, fact checks that did not regard the claim as false, misleading or misinformation were also not further annotated. Fact checking organizations use a variety of heterogeneous labeling systems to indicate veracity. Raters were instructed to err on the side of excluding misinformation claims that were assessed as “possibly true” or “unclear,” etc.
@@ B.1 Image Types
The image type categories were selected to reflect properties of the images that are relevant to an investigator, like a fact checker, engaged in determining its authenticity, as the means used for investigating an image present in a misinformation claim differ based on the nature of the image and the nature of the suspected manipulation. For example, while reverse image search methods are reliable in the case of natural images and are widely deployed in fact checking [18], they may be less useful for images with large amounts of diagrammatic content and text [16], which contain less information than natural images and bear elements that are frequently and identically repeated across images. Additionally, content may be captured by taking a screenshot and re-shared, sometimes several times. Each of these can introduce graphical elements into the new image, and reduce the fidelity of the underlying original.
@@ B.3.1 Fake Official Documents
Fake official documents are created to give the false impression of an official communication from an organization or authority. This category was not present in our original formulation of the typologies, but was added once their prevalence was recognized. These types of images are particularly relevant during public safety events, where communication from officials can be crucial.
@@ Appendix C Data Collection Rater Task
In all cases, raters were instructed to treat the fact check as a source of truth for annotations, and to refrain from making independent judgments. Exceptions to this are cases where the judgement could be made objectively or concerned categories developed specifically for this study (principally, the distinction between basic and complex images, see 5.3) and never when related to the veracity of the misinformation claim being examined or the evidence establishing this veracity. In cases where misinformation claims are addressed by multiple fact checks presenting conflicting assessments of claims, or raters provided conflicting responses, those conflicts will be reflected in the dataset.
@@ C.3 Stage 2: Fine-Grained Annotation of Single-Image Misinformation Claims
Stage 2 contains dozens of categories, most centering on the manipulation present in, or applied to, the image to support or make the misinformation claim (does the image contain AI-based manipulation, did the fact check conduct a reverse-image search, etc). Unlike Stages 1 and 1M, very few annotations were discarded (a total of 69). Each (claim, image) pair has an average of 2.6 Stage 2 annotations, and 84.9% of claims have more than 1 replication.
@@ C.4 Stage 2M: Fine-Grained Annotation of Images from Multiple-Media Misinformation Claims
Stage 2M was the final stage conducted, and is applied to misinformation claims bearing multiple media including at least one image. For every image associated with such a claim, a Stage 2M annotation is applied. Raters are presented with a fact check URL and an original source as identified in Stage 1. Images associated with the claim, identified in Stage 1M (see Sec. C.2), are also shown; one image is the "primary" image, with the remainder secondary. The primary image is the main subject of annotation.